# Efficient Perplexity Bound and Ratio Matching in Discrete Diffusion Language Models

### v1

@@ A.2.1 Rederiving the Denosing Score Entropy Loss
¯$\bar{\ell}\left(\frac{p_{t}({\bm{y}})}{p_{t}({\bm{x}}_{t})},s_{\theta}({\bm{x}}_{t},t)_{\bm{y}}\right),$ (70)
@@
[…]
@@
$=\mathbb{E}_{{\bm{x}}_{t}\sim p_{t}({\bm{x}}_{t})}\bar{\ell}\left(\Sigma_{{\bm{x}}_{0}}\frac{p_{t|0}({\bm{y}}|{\bm{x}}_{0})}{p_{t|0}({\bm{x}}_{t}|{\bm{x}}_{0})}\frac{p_{t|0}({\bm{x}}_{t}|{\bm{x}}_{0})p_{0}({\bm{x}}_{0})}{p_{t}({\bm{x}}_{t})},s_{\theta}({\bm{x}}_{t},t)_{\bm{y}}\right)$ (73)
$=\mathbb{E}_{{\bm{x}}_{t}\sim p_{t}({\bm{x}}_{t})}\bar{\ell}\left(\Sigma_{{\bm{x}}_{0}}\frac{p_{t|0}({\bm{y}}|{\bm{x}}_{0})}{p_{t|0}({\bm{x}}_{t}|{\bm{x}}_{0})}p_{0|t}({\bm{x}}_{0}|{\bm{x}}_{t}),s_{\theta}({\bm{x}}_{t},t)_{\bm{y}}\right)$ (74)
$=\mathbb{E}_{{\bm{x}}_{t}\sim p_{t}({\bm{x}}_{t})}\bar{\ell}\left(\mathbb{E}_{{\bm{x}}_{0}\sim p_{0|t}({\bm{x}}_{0}|{\bm{x}}_{t})}\frac{p_{t|0}({\bm{y}}|{\bm{x}}_{0})}{p_{t|0}({\bm{x}}_{t}|{\bm{x}}_{0})},s_{\theta}({\bm{x}}_{t},t)_{\bm{y}}\right).$ (75)
@@ A.2.2 Deriving Cross-Entropy from Ratio Matching
If we learn the $V$ probabilities $[p^{i}_{0|t}({0}|{\bm{x}}_{t}),p^{i}_{0|t}({1}|{\bm{x}}_{t}),...,p^{i}_{0|t}({(V-1)}|{\bm{x}}_{t})]$, since we analytically have $\frac{p_{t|0}({\bm{y}}^{i}|{h}^{i})}{p_{t|0}({{\bm{x}}_{t}}^{i}|{h}^{i})}$ (from the matrix exponential), then we we will have the ratio $\frac{p_{t}({\bm{y}})}{p_{t}({\bm{x}}_{t})}$. We can choose another sequence ${\bm{z}}$ which also differs only at position $i$ from ${\bm{x}}_{t}$. Then we will have
@@
which highlights the fact that even though the ratios might change, the same $[p^{i}_{0|t}({0}|{\bm{x}}_{t}),p^{i}_{0|t}({1}|{\bm{x}}_{t}),...,p^{i}_{0|t}({(V-1)}|{\bm{x}}_{t})]$ appear as long as the selected position $i$ in the sequence remains unchanged. Therefore if we learn these probabilities for a position, we will have the ratios of all neighbours that differ only in that position, that is, we have $V$ ratios, by modeling these $V$ probabilities.
@@
$-\sum_{i=1}^{L}\mathbb{E}_{t\sim U(0,1)}w(t)\mathbb{E}_{{\bm{x}}_{t}\sim p_{t}({\bm{x}}_{t})}\mathbb{E}_{h\sim p_{0|t}^{i}({h|{\bm{x}}_{t})}}\log{f_{\theta}^{i}({\bm{x}}_{t},t)[h]},$ (89)
@@
$-\sum_{i=1}^{L}\mathbb{E}_{t\sim U(0,1)}\mathbb{E}_{{\bm{x}}_{t}\sim p_{t}({\bm{x}}_{t})}\mathbb{E}_{{\bm{x}}_{0}\sim p_{0|t}({{\bm{x}}_{0}|{\bm{x}}_{t})}}w(t)\log{f_{\theta}^{i}({\bm{x}}_{t},t)[{\bm{x}}_{0}^{i}]}.$ (90)
@@
$=\mathbb{E}_{{\bm{x}}_{0}\sim p_{0}({\bm{x}}_{0})}\mathbb{E}_{{\bm{x}}_{t}\sim p_{t|0}({\bm{x}}_{t}|{\bm{x}}_{0})}\ell\left(\frac{p_{t|0}({\bm{y}}^{i}|{{\bm{x}}_{0}}^{i})}{p_{t|0}({{\bm{x}}_{t}}^{i}|{{\bm{x}}_{0}}^{i})},\Sigma_{h\in[V]}\frac{p_{t|0}({\bm{y}}^{i}|h)}{p_{t|0}({{\bm{x}}_{t}}^{i}|h)}f_{\theta}^{i}({\bm{x}}_{t},t)[h]\right).$ (93)
@@
$-\mathbb{E}_{t\sim U(0,1)}\mathbb{E}_{{\bm{x}}_{0}\sim p_{0}({\bm{x}}_{0})}\mathbb{E}_{{\bm{x}}_{t}\sim p_{t|0}({\bm{x}}_{t}|{\bm{x}}_{0})}\sum_{i=1}^{L}w(t)\log f_{\theta}^{i}({\bm{x}}_{t},t)[{\bm{x}}_{0}^{i}].$ (96)
@@ A.2.3 Relation to Cross Entropy in Continuous Diffusion
[…]
@@
[…]
Clearly, all that need to be learned in order to model the score for the dimensions of the token at position $i$ are the $V$ probabilities $p_{t}({\bm{x}}_{0}^{i}|{\bm{x}}_{t})$ For $L$ such tokens (to model the score of the entire sequence) one simply needs to model $L\times V$ probabilities using cross entropy as in the discrete case.
@@ A.3 Roulette Discrete Diffusion
where $n-1=V$ is the number of tokens in our vocabulary. We add a special token for the absorb (mask) state, therefore increasing the number of total states to $n$, with the token id $n$ corresponding to the absorbed state. Unfortunately, we cannot derive the matrix exponential as in the case of absorb and uniform diffusion, since the corresponding matrix ${\bm{P}}$ of ${\bm{Q}}$ is not idempotent. Therefore, we will first derive the exponential matrix of the uniform and absorb diffusion manually, in order to motivate the manual derivation of the exponential matrix in the roulette case.
@@ A.3.1 Deriving the exponential matrix of the absorb diffusion
If we are at state $i$, the probability of moving at state $n$ (the absorb/mask state) for a time-step of size $\epsilon$ is $\epsilon$. We discretize the time interval into discrete steps that are multiples of $\epsilon$, that is $[0,\tau]\rightarrow{\{\epsilon,...,j\epsilon,...\lfloor\frac{\tau}{\epsilon}\rfloor\epsilon\}}$ and denote the event of jumping from $i\neq n$ to $n$ at time $j\epsilon$ as $a_{j}$. The intersection of any of these two events is empty and their union is the space of all possibilities, thus the probability of being at $n$ at time $\tau$ is
@@ A.3.2 Deriving the exponential of the uniform diffusion
Clearly this is ${\bm{Q}}_{roulette}^{tok}$ when $p_{m}=0$, without the absorption row and column, as our vocabulary size is $V=n-1$ since we do not have the special token. As before we discretize the time interval into discrete steps that are multiples of $\epsilon$, that is $[0,\tau]\rightarrow{\{\epsilon,...,j\epsilon,...\lfloor\frac{\tau}{\epsilon}\rfloor\epsilon\}}$. We assume that initially (at time 00) we start at position $i$. We wish to find the probability of being at state $k\neq i$ at time $\lfloor\frac{\tau}{\epsilon}\rfloor\epsilon$. For the sake of simplicity, we abuse notation by denoting $p_{(\lfloor\frac{\tau}{\epsilon}\rfloor\epsilon-m\epsilon)|0}(j|i)$ as $\bar{p}_{\tau-m\epsilon}(j|i)$. By symmetry we know that for every $j\neq i$ the probability $\bar{p}_{\tau-\epsilon}(j|i)$ is the same. Thus the probability of $\bar{p}_{\tau}(j|i)$ is the sum of the following components:
  • probability of being at $j\not\in\{i,k\}$ at time $\tau-\epsilon$, i.e, ($\bar{p}_{\tau-\epsilon}(j|i)$), times $n-3=V-2$ such states, times probability of moving ($\epsilon$), times probability of hitting $k$ on that move ($\frac{1}{n-1}$).
@@
$+\bar{p}_{\tau-\epsilon}(j|i)(1-\epsilon)+\bar{p}_{\tau-\epsilon}(j|i)\frac{1}{n-1}\epsilon,$ (108)
@@ A.3.3 Deriving the exponential of the roulette diffusion
$+\bar{p}_{\tau-\epsilon}(j|i)(1-\delta)+\bar{p}_{\tau-\epsilon}(j|i)\frac{1}{n-1}\delta$ (116)
@@
$=\frac{(1-p_{m})\epsilon}{(1-p_{m})\epsilon+1(1-\epsilon)}=\alpha\epsilon$ (120)
@@ A.3.5 Time Evolving Roulette
First, we point out that since the size of ${\bm{Q}}_{eroulette}$ is $(n\times n)$ and that of ${\bm{Q}}_{uniform}$ is $(V\times V)$, where $V=n-1$, in order for them to have the same size, we add a row of zeros and a column of zeros to the bottom and right of ${\bm{Q}}_{uniform}$ respectively. Clearly, by definition, these are transition rate matrices, and therefore so is ${\bm{Q}}_{eroulette}^{tok}(t)$. Indeed, the elements in each of its columns add to 00, and the only negative elements are in the diagonals. If we define, $p_{m}(t)$ such that $p_{m}(0)=0$ and $p_{m}(1)=1$, then the limiting distribution will be the one-hot encoding at the absorb state. Furthermore, we can compute the exponential matrix of ${\bm{Q}}_{eroulette}^{tok}(t)$. It is easy to prove that $(1-p_{m}(t))\sigma(t){\bm{Q}}_{uniform}^{tok}$ and $p_{m}(t)\sigma(t){\bm{Q}}_{absorb}^{tok}(t)$ commute with each other, thus
@@
$=e^{p_{m}(t)\sigma_{t}{\bm{Q}}_{absorb}^{tok}}e^{(1-p_{m}(t))\sigma_{t}{\bm{Q}}_{uniform}^{tok}}.$ (130)
@@ B.2 Network Architecture and Hyper-parameters
This core model is grounded in the diffusion transformer architecture introduced by Peebles & Xie (2023), which integrates time conditioning into the conventional encoder-only transformer framework as established by Vaswani et al. (2017); Devlin et al. (2019). However, it includes minor adjustments, such as the use of rotary positional encoding (Su et al., 2024). The model contains approximately 5% more parameters than a standard transformer (utilized in the case of GPT-2), attributed to the incorporation of time conditioning. Additionally, the same tokenizers and data splits as in prior work are employed to avoid introducing artifacts.
@@
In terms of hyperparameters, the model was trained on a single H100 when the sequence length is set at 128, while in the case of sequence lengths of 1024 the model is trained using 8$\times$×H100 with a vocabulary size of 50,257 tokens. Training involves a batch size of 512. The training regime is designed for a total of 400,000 iterations.

### v2

@@ A.2.1 Rederiving the Denosing Score Entropy Loss
$\bar{\ell}\left(\frac{p_{t}({\bm{y}})}{p_{t}({\bm{x}}_{t})},s_{\theta}({\bm{x}}_{t},t)_{\bm{y}}\right),$ (70)
@@
The main bottleneck is that the ratios $\frac{p_{t}({\bm{y}})}{p_{t}({\bm{x}}_{t})}$ are unknown, as $p_{t}({\bm{x}}_{t})=\int p_{t|0}({\bm{x}}_{t}|{\bm{x}}_{0})p_{0}({\bm{x}}_{0})d{\bm{x}}_{0}$ and $p_{0}({\bm{x}}_{0})$ is what we are trying to model in the first place. Luckily, the denoising trick can be employed:
@@
$\mathbb{E}_{{\bm{x}}_{t}\sim p_{t}({\bm{x}}_{t})}\bar{\ell}\left(\Sigma_{{\bm{x}}_{0}}\frac{p_{t|0}({\bm{y}}|{\bm{x}}_{0})}{p_{t|0}({\bm{x}}_{t}|{\bm{x}}_{0})}\frac{p_{t|0}({\bm{x}}_{t}|{\bm{x}}_{0})p_{0}({\bm{x}}_{0})}{p_{t}({\bm{x}}_{t})},s_{\theta}({\bm{x}}_{t},t)_{\bm{y}}\right)$ (73)
$\mathbb{E}_{{\bm{x}}_{t}\sim p_{t}({\bm{x}}_{t})}\bar{\ell}\left(\Sigma_{{\bm{x}}_{0}}\frac{p_{t|0}({\bm{y}}|{\bm{x}}_{0})}{p_{t|0}({\bm{x}}_{t}|{\bm{x}}_{0})}p_{0|t}({\bm{x}}_{0}|{\bm{x}}_{t}),s_{\theta}({\bm{x}}_{t},t)_{\bm{y}}\right)$ (74)
$\mathbb{E}_{{\bm{x}}_{t}\sim p_{t}({\bm{x}}_{t})}\bar{\ell}\left(\mathbb{E}_{{\bm{x}}_{0}\sim p_{0|t}({\bm{x}}_{0}|{\bm{x}}_{t})}\frac{p_{t|0}({\bm{y}}|{\bm{x}}_{0})}{p_{t|0}({\bm{x}}_{t}|{\bm{x}}_{0})},s_{\theta}({\bm{x}}_{t},t)_{\bm{y}}\right).$ (75)
@@ A.2.2 Deriving Cross-Entropy from Ratio Matching
If we learn the $V$ probabilities $p^{i}_{0|t}({0}|{\bm{x}}_{t}),p^{i}_{0|t}({1}|{\bm{x}}_{t}),...,p^{i}_{0|t}({(V-1)}|{\bm{x}}_{t})]$, since we analytically have $\frac{p_{t|0}({\bm{y}}^{i}|{h}^{i})}{p_{t|0}({{\bm{x}}_{t}}^{i}|{h}^{i})}$ (from the matrix exponential), then we we will have the ratio $\frac{p_{t}({\bm{y}})}{p_{t}({\bm{x}}_{t})}$. We can choose another sequence $\bm{z}$} which also differs only at position $i$ from ${\bm{x}}_{t}$. Then we will have
@@
which highlights the fact that even though the ratios might change, the same $p^{i}_{0|t}({0}|{\bm{x}}_{t}),p^{i}_{0|t}({1}|{\bm{x}}_{t}),...,p^{i}_{0|t}({(V-1)}|{\bm{x}}_{t})]$ appear as long as the selected position $i$ in the sequence remains unchanged. Therefore if we learn these probabilities for a position, we will have the ratios of all neighbours that differ only in that position, that is, we have $V$ ratios, by modeling these $V$ probabilities.
@@
$\sum_{i=1}^{L}\mathbb{E}_{t\sim U(0,1)}w(t)\mathbb{E}_{{\bm{x}}_{t}\sim p_{t}({\bm{x}}_{t})}\mathbb{E}_{h\sim p_{0|t}^{i}({h|{\bm{x}}_{t})}}\log{f_{\theta}^{i}({\bm{x}}_{t},t)[h]},$ (89)
@@
$\sum_{i=1}^{L}\mathbb{E}_{t\sim U(0,1)}\mathbb{E}_{{\bm{x}}_{t}\sim p_{t}({\bm{x}}_{t})}\mathbb{E}_{{\bm{x}}_{0}\sim p_{0|t}({{\bm{x}}_{0}|{\bm{x}}_{t})}}w(t)\log{f_{\theta}^{i}({\bm{x}}_{t},t)[{\bm{x}}_{0}^{i}]}.$ (90)
@@
$\mathbb{E}_{{\bm{x}}_{0}\sim p_{0}({\bm{x}}_{0})}\mathbb{E}_{{\bm{x}}_{t}\sim p_{t|0}({\bm{x}}_{t}|{\bm{x}}_{0})}\ell\left(\frac{p_{t|0}({\bm{y}}^{i}|{{\bm{x}}_{0}}^{i})}{p_{t|0}({{\bm{x}}_{t}}^{i}|{{\bm{x}}_{0}}^{i})},\Sigma_{h\in[V]}\frac{p_{t|0}({\bm{y}}^{i}|h)}{p_{t|0}({{\bm{x}}_{t}}^{i}|h)}f_{\theta}^{i}({\bm{x}}_{t},t)[h]\right).$ (93)
@@
$\mathbb{E}_{t\sim U(0,1)}\mathbb{E}_{{\bm{x}}_{0}\sim p_{0}({\bm{x}}_{0})}\mathbb{E}_{{\bm{x}}_{t}\sim p_{t|0}({\bm{x}}_{t}|{\bm{x}}_{0})}\sum_{i=1}^{L}w(t)\log f_{\theta}^{i}({\bm{x}}_{t},t)[{\bm{x}}_{0}^{i}].$ (96)
@@ A.2.3 Relation to Cross Entropy in Continuous Diffusion
$\nabla_{{\bm{x}}_{t}^{i}}\log p_{t}({\bm{x}}_{t})=\int\frac{{\bm{x}}_{0}^{i}}{\sigma_{t}^{2}}p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})d{\bm{x}}_{0}-\frac{{\bm{x}}_{t}^{i}}{\sigma_{t}^{2}}=\int\frac{{\bm{x}}_{0}^{i}}{\sigma_{t}^{2}}\left(\int p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})d{\bm{x}}_{0}^{-i}\right)d{\bm{x}}_{0}^{i}-\frac{{\bm{x}}_{t}^{i}}{\sigma_{t}^{2}},$ (99)
@@
$s_{\theta}({\bm{x}}_{t},t)_{(i1,i2,..,iD)}\approx\nabla_{{\bm{x}}_{t}^{i}}\log p_{t}({\bm{x}}_{t})=\int\frac{{\bm{x}}_{0}^{i}}{\sigma_{t}^{2}}p_{t}({\bm{x}}_{0}^{i}|{\bm{x}}_{t})d{\bm{x}}_{0}^{i}-\frac{{\bm{x}}_{t}^{i}}{\sigma_{t}^{2}}.$ (100)
Clearly, all that need to be learned in order to model the score for the dimensions of the token at position $i$ are the $V$ probabilities $p_{t}({\bm{x}}_{0}^{i}|{\bm{x}}_{t})$ . For $L$ such tokens (to model the score of the entire sequence) one simply needs to model $L\times V$ probabilities using cross entropy as in the discrete case.
@@ A.3 Roulette Discrete Diffusion
where n1=Vn-1=V is the number of tokens in our vocabulary. We add a special token for the absorb (mask) state, therefore increasing the number of total states to $n$, with the token id $n$ corresponding to the absorbed state. Unfortunately, we cannot derive the matrix exponential as in the case of absorb and uniform diffusion, since the corresponding matrix $\bm{P}$} of $\bm{Q}$} is not idempotent. Therefore, we will first derive the exponential matrix of the uniform and absorb diffusion manually, in order to motivate the manual derivation of the exponential matrix in the roulette case.
@@ A.3.1 Deriving the exponential matrix of the absorb diffusion
If we are at state $i$, the probability of moving at state $n$ (the absorb/mask state) for a time-step of size $\epsilon$ is $\epsilon$. We discretize the time interval into discrete steps that are multiples of $\epsilon$, that is $0,\tau]\rightarrow{\{\epsilon,...,j\epsilon,...\lfloor\frac{\tau}{\epsilon}\rfloor\epsilon\}}$ and denote the event of jumping from $i\neq n$ to $n$ at time $j\epsilon$ as $a_{j}$. The intersection of any of these two events is empty and their union is the space of all possibilities, thus the probability of being at $n$ at time $\tau$ is
@@ A.3.2 Deriving the exponential of the uniform diffusion
Clearly this is ${\bm{Q}}_{roulette}^{tok}$ when $p_{m}=0$, without the absorption row and column, as our vocabulary size is V=n1V=n-1 since we do not have the special token. As before we discretize the time interval into discrete steps that are multiples of $\epsilon$, that is $0,\tau]\rightarrow{\{\epsilon,...,j\epsilon,...\lfloor\frac{\tau}{\epsilon}\rfloor\epsilon\}}$. We assume that initially (at time 0) we start at position $i$. We wish to find the probability of being at state $k\neq i$ at time $\lfloor\frac{\tau}{\epsilon}\rfloor\epsilon$. For the sake of simplicity, we abuse notation by denoting $p_{(\lfloor\frac{\tau}{\epsilon}\rfloor\epsilon-m\epsilon)|0}(j|i)$ as $\bar{p}_{\tau-m\epsilon}(j|i)$. By symmetry we know that for every $j\neq i$ the probability $\bar{p}_{\tau-\epsilon}(j|i)$ is the same. Thus the probability of $\bar{p}_{\tau}(j|i)$ is the sum of the following components:
  • probability of being at $j\not\in\{i,k\}$ at time $\tau-\epsilon$, i.e, ($\bar{p}_{\tau-\epsilon}(j|i)$), times n3=V2n-3=V-2 such states, times probability of moving ($\epsilon$), times probability of hitting $k$ on that move ($\frac{1}{n-1}$).
@@
$\bar{p}_{\tau-\epsilon}(j|i)(1-\epsilon)+\bar{p}_{\tau-\epsilon}(j|i)\frac{1}{n-1}\epsilon,$ (108)
@@ A.3.3 Deriving the exponential of the roulette diffusion
$\bar{p}_{\tau-\epsilon}(j|i)(1-\delta)+\bar{p}_{\tau-\epsilon}(j|i)\frac{1}{n-1}\delta$ (116)
@@
$\frac{(1-p_{m})\epsilon}{(1-p_{m})\epsilon+1(1-\epsilon)}=\alpha\epsilon$ (120)
@@ A.3.5 Time Evolving Roulette
First, we point out that since the size of ${\bm{Q}}_{eroulette}$ is $(n\times n)$ and that of ${\bm{Q}}_{uniform}$ is $(V\times V)$, where V=n1V=n-1, in order for them to have the same size, we add a row of zeros and a column of zeros to the bottom and right of ${\bm{Q}}_{uniform}$ respectively. Clearly, by definition, these are transition rate matrices, and therefore so is ${\bm{Q}}_{eroulette}^{tok}(t)$. Indeed, the elements in each of its columns add to 0, and the only negative elements are in the diagonals. If we define, $p_{m}(t)$ such that $p_{m}(0)=0$ and $p_{m}(1)=1$, then the limiting distribution will be the one-hot encoding at the absorb state. Furthermore, we can compute the exponential matrix of ${\bm{Q}}_{eroulette}^{tok}(t)$. It is easy to prove that $(1-p_{m}(t))\sigma(t){\bm{Q}}_{uniform}^{tok}$ and $p_{m}(t)\sigma(t){\bm{Q}}_{absorb}^{tok}(t)$ commute with each other, thus
@@
$e^{p_{m}(t)\sigma_{t}{\bm{Q}}_{absorb}^{tok}}e^{(1-p_{m}(t))\sigma_{t}{\bm{Q}}_{uniform}^{tok}}.$ (130)
@@ B.2 Network Architecture and Hyper-parameters
This core model is grounded in the diffusion transformer architecture introduced by Peebles and Xie (2023), which integrates time conditioning into the conventional encoder-only transformer framework as established by Vaswani et al. (2017); Devlin et al. (2019). However, it includes minor adjustments, such as the use of rotary positional encoding (Su et al., 2024). The model contains approximately 5% more parameters than a standard transformer (utilized in the case of GPT-2), attributed to the incorporation of time conditioning. Additionally, the same tokenizers and data splits as in prior work are employed to avoid introducing artifacts.
@@
[…]